OpenStudy (anonymous):

log(base3)(x-4)+log(base2)(x+4)=2 I don't know how to solve this when the logs have different bases. As I understand they have to be the same?

OpenStudy (jtvatsim):

have you been taught the base conversion formula? \[\log_a{x} = \frac{\ln{x}}{\ln{a}}\]

OpenStudy (anonymous):

Yes very vaguely though. We are behind in the semester so we rushed through this section and I have 10 pages of notes but no clue what to do

OpenStudy (jtvatsim):

ok, i'm still working on the problem myself... it's not obvious. :)

OpenStudy (anonymous):

okay I will tell you what I think it might be too

OpenStudy (anonymous):

ya I have no IDEA at all.... I was thinking it could work with the condensing formula but thats not the case

OpenStudy (jtvatsim):

there's another conversion formula for any base I'm looking into that right now

OpenStudy (jtvatsim):

This site has some good examples: http://www.priklady.eu/en/Mathematics/Logarithmic-Equations/Logarithmic-Equations-Other-Bases.alej

OpenStudy (anonymous):

is that a foreign language?? Lol I have no idea how to convert that to my problem.

OpenStudy (jtvatsim):

maybe, I just click on the show solution link below each question to see the answers steps

OpenStudy (anonymous):

okay, I will look more into that question. I have a few others Im going to post.

OpenStudy (jtvatsim):

kk